OpenStudy (anonymous):

write the single logarithm 4 log x-6 log (x+2)

OpenStudy (anonymous):

i got log x^4(x+2)^6 not sure if right

OpenStudy (anonymous):

log(a)-log(b)=log(a/b)

OpenStudy (anonymous):

so what would you do after you get log(x^4)-log(x+2)^6

OpenStudy (anonymous):

noy real sure kinda guessing

OpenStudy (anonymous):

its a law of logarithms \[\log a- \log b=\log (a/b)\]

OpenStudy (anonymous):

log x(x+2)^24

OpenStudy (anonymous):

Divide them.

OpenStudy (anonymous):

you mean like 24 log x/x+2

OpenStudy (anonymous):

No i mean what we got earlier log(x^4)-log(x+2)^6

OpenStudy (anonymous):

not getting it thanks anyways

OpenStudy (anonymous):

Its kinda like the other one you did just with a different law.... instead of log(a)+log(b)=log(a*b) its log(a)-log(b)=log(a/b)

OpenStudy (anonymous):

4 logx- 6/log(x+2) maybe

OpenStudy (anonymous):

\[\huge \log (x^4/(x-6)^{2})\]

OpenStudy (anonymous):

so thats how you write it as a single logarithm

OpenStudy (anonymous):

Yes. Here's the laws of logarithms you should get familiar with

OpenStudy (anonymous):

ok thanks for the help